# Top-quark pair production as a probe of light top-philic scalars and anomalous Higgs interactions

### v1

@@ 1 Introduction
On the theoretical side, substantial advancements have been made over the past fifteen years. Achieving next-to-next-to-leading order (NNLO) precision in Quantum Chromodynamics (QCD) Baernreuther:2012ws ; Czakon:2012zr ; Czakon:2012pz ; Czakon:2013goa ; Czakon:2015owf ; Czakon:2016dgf ; Catani:2019iny ; Catani:2019hip ; Catani:2020tko alongside the calculation and the combination with the of electroweak (EW) corrections  Hollik:2011ps ; Kuhn:2011ri ; Bernreuther:2012sx ; Pagani:2016caq ; Czakon:2017wor ; Czakon:2017lgo ; Gutschow:2018tuk ; Frederix:2018nkq ; Czakon:2019txp has been a significant milestone, with various independent collaborations attaining this level of accuracy at a fully differential level. Moreover, the achievement of fully exclusive predictions, which can be seamlessly integrated into simulations, has pushed the boundaries to NNLO precision coupled with parton shower (PS) accuracy Mazzitelli:2020jio ; Mazzitelli:2021mmm . Advancements have been made also in the calculation final state signatures. In certain regions of phase space, processes that do not strictly feature one or two on-shell top quarks have been computed with next-to-leading order (NLO) precision in QCD and EW Denner:2016jyo ; Denner:2017kzu ; Denner:2023grl and in some cases even coupled with parton showers Jezo:2016ujg ; Jezo:2023rht , broadening the scope of the analyses. Higher order corrections, up to NNLO in QCD, have also been calculated for spin-correlations Czakon:2020qbd ; Frederix:2021zsh . Resummation effects, crucial particularly at threshold, have also been included up to next-to-next-to-leading logarithmic (NNLL) accuracy Ahrens:2010zv ; Ferroglia:2012ku ; Pecjak:2016nee ; Czakon:2019txp . Looking ahead, efforts towards achieving next-to-next-to-next-to-leading order ($\rm N^{3}LO$) QCD accuracy have started, with the calculation for the top-quark decay that has recently appeared in the literature Chen:2023osm . With the extended runtime of the High-Luminosity LHC (HL-LHC) in mind, such a precision appears increasingly attainable.
Given what we have already at disposal and what is expected in the near and medium term is therefore mandatory to think how to best leverage such a phenomenal combination of theoretical predictions and experimental data. Fortunately, the top quark, occupying a unique position within the Standard Model as both the sole naturally occurring quark with a mass on the order of the electroweak scale, offers an ideal platform for probing novel concepts aimed at unraveling the mysteries surrounding electroweak symmetry breaking and investigating the potential existence of new phenomena.
@@
The paper is organised as follows. Sections 24 concern the case of the light top-philic scalar $S$, while Sections 57 the case of the Higgs boson. In Sec. 2 we present the theoretical framework and we give the details of the calculation of virtual effects from a top-philic scalar on the $t\bar{t}$ cross section. We also show how to combine these effects with precise SM predictions. In Sec. 3 we discuss the different pattern that can arise depending on how is the interaction between the top-quark and the scalar $S$, namely, if it is purely CP-even, CP-odd or a mixture of the two. We then show in Sec. 4 the sensitivity that we can already achieve with current data on such interactions. In Sec. 5 we show how the calculation for the scalar $S$ presented in Sec. 2 can be recycled for the case of the Higgs boson and we also discuss the connection with the SMEFT framework. We then show the impact of the virtual effects from the Higgs on the $t\bar{t}$ cross section and we discuss in details the differences with the calculation in Ref. Martini:2021uey . In Sec. 7 we discuss the bounds that can be set on CP-even and CP-odd Yukawa interactions of the top-quark. Finally, we give our conclusions in Sec. 8. In the Appendices we collect useful analytical formulas which we used for cross-checking our numerical simulations and additional information entering the statistical analysis.
@@ 2.1 Lagrangian and notation
where $|C_{t}|$ is parametrising the “total” strength of the interaction between the top quark and and the scalar $S$. Instead, the admixture of the CP-even and CP-odd components is parameterised by the phase $\phi$.
@@ 2.2 Scalar one-loop corrections to top-quark pair production
which is the QCD Lagrangian with only one fermion, the top quark. This means that the only interactions that can be renormalised is the one between the gluons and the top quark.
Second, according to Eq. (8), one-loop corrections can in principle induce effects of order $c_{t}^{2}$, $\tilde{c}_{t}^{2}$ and also $c_{t}\tilde{c}_{t}$. This means that also the counterterms can include these same three classes of corrections, such that all UV divergencies are canceled.
@@
with $\widehat{\Sigma}^{1}_{\rm NP}$ being the UV divergent one-loop corrections (the associated diagram is shown in Fig. 17 in Appendix A, where we also calculate it) while the remaining contributions originates from the associated UV counterterm. The on-shell scheme corresponds to requiring
@@
Because of the optical theorem, $2\to 2$ tree-level amplitudes as $\mathcal{M}^{0}_{\rm SM}$ have to be real and therefore any contribution that has an imaginary component in $\mathcal{M}^{1}_{\rm NP}$ is not entering $\sigma_{\rm NP}$. In other words, even if $\Sigma_{P}$ is divergent and not renormalised, it is not entering our results, similarly to any other possible imaginary component from loop diagrams in $\mathcal{M}^{1}_{\rm NP}$. We will understood this technical detail in the following, but it is important to remember that for a generic calculation an additional counterterm $\delta_{\tilde{m}_{t}}$ is necessary and cannot be ignored. The presence of the subscript QCD in $\sigma_{\rm LO_{\rm QCD}}$ and $\sigma_{\rm LO_{QCD}+NP}$ is a reminder that at the tree level additional perturbative orders are possible; this will be manifest in Sec. 2.3.
Once the quantities $\Sigma_{V}$ and $\Sigma_{S}$ are known, it is possible to calculate $\mathcal{M}^{1}_{\rm NP}$ for both the $q\bar{q}$ and $gg$ initiated processes. The unrenormalised one-loop amplitude $\widehat{{\mathcal{M}}^{1}_{\rm NP}}$ corresponds to the calculation of the one-loop diagrams in Fig. 1, excluding those with a loop on the external top-quark legs. Following the LSZ theorem for scattering amplitudes, in the on-shell scheme the associated contribution is exactly canceled by the $\sqrt{Z_{\psi_{t}}}$ normalisation of the external field. The amplitude associated with the UV counter terms $\mathcal{M}^{1}_{\rm NP,\,CT}$ corresponds to the diagrams in Fig. 4, where the Feynman rules are specified in Fig. 3 and in turn by Eqs. (27) and (28).
@@
We emphasise that such UFO model is very similar to the one used to perform the study of top-quark pair production of Ref. Blasi:2023hvb , where instead of a generic scalar $S$ a top-philic ALP $a$ has been considered. The main differences are two. First, in the case of the ALP an additional interaction between $a$ and the gluons is present, unlike the case of $S$. The technical details regarding this additional interaction have been described in depth in Ref. Blasi:2023hvb , especially Appendices C and D of that work. Second, $S$ can feature both CP-odd and CP-even interactions, violating the CP symmetry, while the top-philic ALP in the model of Ref. Blasi:2023hvb is itself CP-odd and features only CP conserving interactions, namely $c_{t}=0$ in the notation of this paper. Excluding the trivial interaction between $a$ and the gluons, the calculation presented in this work is therefore a generalisation of the one presented in Ref. Blasi:2023hvb and we provide in this work, especially in this section and in Appendices A and B, further technical details that were not discussed in Ref. Blasi:2023hvb . For the purely $\tilde{c}_{t}$-dependent component and excluding the interaction between $a$ and the gluons, our calculation is equivalent not only to what has been presented in Ref. Blasi:2023hvb , but also to the results for the $gg\to t\bar{t}$ process in Ref. Phan:2023dqw . [^3] In Ref. Phan:2023dqw , however, the $\overline{\rm MS}$ scheme has been used for the renormalisation of the top-quark wave function. This difference have anyway an impact only on Green functions and not on scattering amplitudes, but only the latter are relevant in our study.
@@
where the dependence on $c_{t}$ and $\tilde{c}_{t}$ has been made explicit. It is easy to see the contribution proportional to both $c_{t}$ and $\tilde{c}_{t}$
@@
This is true not only for the virtual corrections but also in the case of the emission of a real scalar $S$. This property is consistent also with the fact $\Sigma_{P}$ proportional to $c_{t}\tilde{c}_{t}$ and divergent, yet as explained before it does not contribute to $\sigma_{\rm NP}$. It is also important to note that the quantities $\bar{\sigma}_{c_{t}}$ and $\bar{\sigma}_{\tilde{c}_{t}}$, which at this stage of the discussion involve only virtual corrections, can also be negative and there can be also large cancellations among them.
@@ 3 Scalar $S$: Numerical results and phenomenology
Given the dependence of $\sigma_{\rm NP}$ on only $c_{t}^{2}$ and $\tilde{c}_{t}^{2}$, with no mixed $c_{t}\tilde{c}_{t}$ terms, the first two cases are sufficient for extrapolating the size of $\sigma_{\rm NP}$ in any configuration; they correspond to $\bar{\sigma}_{c_{t}}$ and $\bar{\sigma}_{\tilde{c}_{t}}$ in Eq. (31), respectively. The third case is useful to assess presence of possible cancellations, which as we will see, happen at the threshold where $\bar{\sigma}_{c_{t}}$ is positive while $\bar{\sigma}_{\tilde{c}_{t}}$ is negative.
@@
The shapes and especially the sign of the $\sigma_{\rm NP}/\sigma_{\rm LO_{\rm QCD}}$ ratio is very different for the $m(t\bar{t})$ distributions of the CP-even and CP-odd cases. In the purely scalar case, corrections are negative and of the order $-c_{t}^{2}\times 10\%$ for large invariant masses for $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$. For larger values of $m_{S}$, corrections are still negative but smaller and smaller in absolute value by increasing $m_{S}$. Going towards the threshold region instead, they reach values a bit smaller than $+c_{t}^{2}\times 10\%$ for $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$ and instead smaller for lager $m_{S}$ values. In the purely pseudoscalar case, on the contrary, corrections are always negative. For $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$, corrections are of the order of $-c_{t}^{2}\times 2\%$ for large invariant masses, growing in absolute value up to $-c_{t}^{2}\times 5\%$ at the threshold. Increasing $m_{S}$ the absolute value of the corrections decreases for large $m(t\bar{t})$ (halved for $m_{S}=300\leavevmode\nobreak\ {\rm GeV}$) while it grows at the threshold (doubled for $m_{S}=300\leavevmode\nobreak\ {\rm GeV}$). As already said, the maximally mixed case (right plot) shows distributions that correspond to the average of the analogous ones in the left and central plot. Clearly, by varying the angle $\phi$ that parameterises the mixing of the CP-even and CP-odd component, very different shapes can be obtained.
@@ 4.1 Statistical method, theory predictions and data
Our goal is to present an exploratory study for the potential bounds that can be obtained at the LHC on the $c_{t}$ (CP-even) and $\tilde{c}_{t}$ (CP-odd) couplings. We neither aim to fully exploit the potential of differential distributions that can be already now measured with high accuracy at the LHC, nor to quantify the projections for the full data-set of the High-Luminosity program; these aspects are left for future studies. Here, we want simply to show how the measurements of such process can lead to constrains for $c_{t}$ (CP-even) and $\tilde{c}_{t}$ (CP-odd) couplings of a top-philic additional scalar and the estimate the corresponding bounds.
@@
The theory prediction $\vec{\sigma}_{\rm th.}$ is given by $\sigma^{\rm mult.}_{\rm SM+NP}$, see the expression in Eq. (39), evaluated for the same binning of the data considered, $\vec{\sigma}_{\mathrm{obs.}}$. In fact, for our exploratory study we use pseudo-data for $\vec{\sigma}_{\mathrm{obs.}}$. First, we have considered several top-quark pair analyses ATLAS:2015lsn ; ATLAS:2016pal ; ATLAS:2019hxz ; CMS:2016oae ; CMS:2018adi ; CMS:2018htd in order to select a binning that is not only useful for our fit but also realistic. Given features discussed in Sec. 3, we have identified the CMS measurement in the lepton+jets channel of Ref. CMS:2018htd for the $m(t\bar{t})$ distribution as the most useful, since it has a large number of bins in the low $m(t\bar{t}$) region. Then, we have generated pseudo data assuming the SM only [^8] In this work, the only exception to this procedure is what is done in Sec. 7.2, as explained therein.: we have calculated for the binning of $m(t\bar{t})$ of Ref. CMS:2018htd $\sigma^{\rm mult.}_{\rm SM+NP}$ with $c_{t}=\tilde{c}_{t}=0$, in other words, the best SM best prediction, the quantity $\sigma^{\rm mult.}_{\rm SM}$ defined in Eq. (36). This procedure fully defines the quantities $\vec{\sigma}_{\mathrm{obs.}}$ and $\vec{\sigma}_{\rm th.}$. Further information entering the fit (binning, values of $\sigma^{\rm mult.}_{\rm SM+NP}$, etc.) is reported in Appendix. D.
The last piece of information that has to be specified is the covariance matrix $\mathbf{V}^{-1}$. In Sec. 7 we will perform the same study discussed in this Section with the Higgs boson $H$ in the place of $S$. For consistency, $\vec{\sigma}_{\mathrm{obs.}}$ and $\mathbf{V}^{-1}$ will be the same used here and in particular it will be more clear the choice we have taken for $\mathbf{V}^{-1}$ and we explain in the following. Starting from the value of $\mathbf{V}^{-1}$ in Ref. CMS:2018htd , we have rescaled the entries of the matrix by a factor $2.5$, such that in the case of the Higgs boson discussed in Sec. 7 we find for the CP-even coupling of top-Higgs coupling the same uncertainty reported in the aforecited Refs. CMS:2019art ; CMS:2020djy . In this way we can mimic, in a simplified statistical framework, the precision already achieved in the experimental analyses based on an integrated luminosity of $35.8\,\rm fb^{-1}$.

### v2

@@ 1 Introduction
On the theoretical side, substantial advancements have been made over the past fifteen years. Achieving next-to-next-to-leading order (NNLO) precision in Quantum Chromodynamics (QCD) Baernreuther:2012ws ; Czakon:2012zr ; Czakon:2012pz ; Czakon:2013goa ; Czakon:2015owf ; Czakon:2016dgf ; Catani:2019iny ; Catani:2019hip ; Catani:2020tko alongside the calculation and the combination with the electroweak (EW) corrections  Hollik:2011ps ; Kuhn:2011ri ; Bernreuther:2012sx ; Pagani:2016caq ; Czakon:2017wor ; Czakon:2017lgo ; Gutschow:2018tuk ; Frederix:2018nkq ; Czakon:2019txp has been a significant milestone, with various independent collaborations attaining this level of accuracy at a fully differential level. Moreover, the achievement of fully exclusive predictions, which can be seamlessly integrated into simulations, has pushed the boundaries to NNLO precision coupled with parton shower (PS) accuracy Mazzitelli:2020jio ; Mazzitelli:2021mmm . Advancements have been made also in the calculation of final state signatures. In certain regions of phase space, processes that do not strictly feature one or two on-shell top quarks have been computed with next-to-leading order (NLO) precision in QCD and EW Denner:2016jyo ; Denner:2017kzu ; Denner:2023grl and in some cases even coupled with parton showers Jezo:2016ujg ; Jezo:2023rht , broadening the scope of the analyses. Higher order corrections, up to NNLO in QCD, have also been calculated for spin-correlations Czakon:2020qbd ; Frederix:2021zsh . Resummation effects, crucial particularly at threshold, have also been included up to next-to-next-to-leading logarithmic (NNLL) accuracy Ahrens:2010zv ; Ferroglia:2012ku ; Pecjak:2016nee ; Czakon:2019txp . Looking ahead, efforts towards achieving next-to-next-to-next-to-leading order ($\rm N^{3}LO$) QCD accuracy have started, with the calculation for the top-quark decay that has recently appeared in the literature Chen:2023osm . With the extended runtime of the High-Luminosity LHC (HL-LHC) in mind, such a precision appears increasingly attainable.
Given what we have already at disposal and what is expected in the near and medium term, it is therefore mandatory to think how to best leverage such a phenomenal combination of theoretical predictions and experimental data. Fortunately, the top quark, occupying a unique position within the Standard Model as both the sole naturally occurring quark with a mass on the order of the electroweak scale, offers an ideal platform for probing novel concepts aimed at unravelling the mysteries surrounding electroweak symmetry breaking and investigating the potential existence of new phenomena.
@@
The paper is organised as follows. Sections 24 concern the case of the light top-philic scalar $S$, while Sections 57 the case of the Higgs boson. In Sec. 2 we present the theoretical framework and we give the details of the calculation of virtual effects from a top-philic scalar on the $t\bar{t}$ cross section. We also show how to combine these effects with precise SM predictions. In Sec. 3 we discuss the different patterns that can arise depending on how is the interaction between the top-quark and the scalar $S$, namely, if it is purely CP-even, CP-odd or a mixture of the two. We then show in Sec. 4 the sensitivity that we can already achieve with current data on such interactions. In Sec. 5 we show how the calculation for the scalar $S$ presented in Sec. 2 can be recycled for the case of the Higgs boson and we also discuss the connection with the SMEFT framework. We then show the impact of the virtual effects from the Higgs on the $t\bar{t}$ cross section and we discuss in details the differences with the calculation in Ref. Martini:2021uey . In Sec. 7 we discuss the bounds that can be set on CP-even and CP-odd Yukawa interactions of the top-quark. Finally, we give our conclusions in Sec. 8. In the Appendices we collect useful analytical formulas which we used for cross-checking our numerical simulations and additional information entering the statistical analysis.
@@ 2.1 Lagrangian and notation
where $|C_{t}|$ is parametrising the “total” strength of the interaction between the top quark and the scalar $S$. Instead, the admixture of the CP-even and CP-odd components is parameterised by the phase $\phi$.
@@ 2.2 Scalar one-loop corrections to top-quark pair production
which is the QCD Lagrangian with only one fermion, the top quark. This means that the only interaction that can be renormalised is the one between the gluons and the top quark.
Second, according to Eq. (8), one-loop corrections can in principle induce effects of order $c_{t}^{2}$, $\tilde{c}_{t}^{2}$ and also $c_{t}\tilde{c}_{t}$. This means that also the counterterms can include these same three classes of corrections, such that all UV divergencies are cancelled.
@@
with $\widehat{\Sigma}^{1}_{\rm NP}$ being the UV divergent one-loop corrections (the associated diagram is shown in Fig. 17 in Appendix A, where we also calculate it) while the remaining contributions originate from the associated UV counterterm. The on-shell scheme corresponds to requiring
@@
Because of the optical theorem, $2\to 2$ tree-level amplitudes as $\mathcal{M}^{0}_{\rm SM}$ have to be real and therefore any contribution that has an imaginary component in $\mathcal{M}^{1}_{\rm NP}$ is not entering $\sigma_{\rm NP}$. In other words, even if $\Sigma_{P}$ is divergent and not renormalised, it is not entering our results, similarly to any other possible imaginary component from loop diagrams in $\mathcal{M}^{1}_{\rm NP}$. We will understand this technical detail in the following, but it is important to remember that for a generic calculation an additional counterterm $\delta_{\tilde{m}_{t}}$ is necessary and cannot be ignored. The presence of the subscript QCD in $\sigma_{\rm LO_{\rm QCD}}$ and $\sigma_{\rm LO_{QCD}+NP}$ is a reminder that at the tree level additional perturbative orders are possible; this will be manifest in Sec. 2.3.
Once the quantities $\Sigma_{V}$ and $\Sigma_{S}$ are known, it is possible to calculate $\mathcal{M}^{1}_{\rm NP}$ for both the $q\bar{q}$ and $gg$ initiated processes. The unrenormalised one-loop amplitude $\widehat{{\mathcal{M}}^{1}_{\rm NP}}$ corresponds to the calculation of the one-loop diagrams in Fig. 1, excluding those with a loop on the external top-quark legs. Following the LSZ theorem for scattering amplitudes, in the on-shell scheme the associated contribution is exactly cancelled by the $\sqrt{Z_{\psi_{t}}}$ normalisation of the external field. The amplitude associated with the UV counter terms $\mathcal{M}^{1}_{\rm NP,\,CT}$ corresponds to the diagrams in Fig. 4, where the Feynman rules are specified in Fig. 3 and in turn by Eqs. (27) and (28).
@@
We emphasise that such UFO model is very similar to the one used to perform the study of top-quark pair production of Ref. Blasi:2023hvb , where instead of a generic scalar $S$ a top-philic ALP $a$ has been considered. The main differences are two. First, in the case of the ALP an additional interaction between $a$ and the gluons is present, unlike the case of $S$. The technical details regarding this additional interaction have been described in depth in Ref. Blasi:2023hvb , especially Appendices C and D of that work. Second, $S$ can feature both CP-odd and CP-even interactions, violating the CP symmetry, while the top-philic ALP in the model of Ref. Blasi:2023hvb is itself CP-odd and features only CP conserving interactions, namely $c_{t}=0$ in the notation of this paper. Excluding the trivial interaction between $a$ and the gluons, the calculation presented in this work is therefore a generalisation of the one presented in Ref. Blasi:2023hvb and we provide in this work, especially in this section and in Appendices A and B, further technical details that were not discussed in Ref. Blasi:2023hvb . For the purely $\tilde{c}_{t}$-dependent component and excluding the interaction between $a$ and the gluons, our calculation is equivalent not only to what has been presented in Ref. Blasi:2023hvb , but also to the results for the $gg\to t\bar{t}$ process in Ref. Phan:2023dqw . [^3] In Ref. Phan:2023dqw , however, the $\overline{\rm MS}$ scheme has been used for the renormalisation of the top-quark wave function. This difference has anyway an impact only on Green functions and not on scattering amplitudes, but only the latter are relevant in our study.
@@
where the dependence on $c_{t}$ and $\tilde{c}_{t}$ has been made explicit. It is easy to see that the contribution proportional to both $c_{t}$ and $\tilde{c}_{t}$ vanishes, i.e.
@@
This is true not only for the virtual corrections but also in the case of the emission of a real scalar $S$. This property is also consistent with the fact that $\Sigma_{P}$ is proportional to $c_{t}\tilde{c}_{t}$ and it is divergent. Indeed, as explained before $\Sigma_{P}$ anyway does not contribute to $\sigma_{\rm NP}$. It is also important to note that the quantities $\bar{\sigma}_{c_{t}}$ and $\bar{\sigma}_{\tilde{c}_{t}}$, which at this stage of the discussion involve only virtual corrections, can also be negative and there can be also large cancellations among them.
@@ 3 Scalar $S$: Numerical results and phenomenology
Given the dependence of $\sigma_{\rm NP}$ on only $c_{t}^{2}$ and $\tilde{c}_{t}^{2}$, with no mixed $c_{t}\tilde{c}_{t}$ terms, the first two cases are sufficient for extrapolating the size of $\sigma_{\rm NP}$ in any configuration; they correspond to $\bar{\sigma}_{c_{t}}$ and $\bar{\sigma}_{\tilde{c}_{t}}$ in Eq. (31), respectively. The third case is useful to assess the presence of possible cancellations, which as we will see, happen at the threshold where $\bar{\sigma}_{c_{t}}$ is positive while $\bar{\sigma}_{\tilde{c}_{t}}$ is negative.
@@
The shapes and especially the sign of the $\sigma_{\rm NP}/\sigma_{\rm LO_{\rm QCD}}$ ratio are very different for the $m(t\bar{t})$ distributions of the CP-even and CP-odd cases. In the purely scalar case, corrections are negative and of the order $-c_{t}^{2}\times 10\%$ for large invariant masses for $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$. For larger values of $m_{S}$, corrections are still negative but smaller and smaller in absolute value by increasing $m_{S}$. Going towards the threshold region instead, they reach values a bit smaller than $+c_{t}^{2}\times 10\%$ for $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$ and instead smaller for lager $m_{S}$ values. In the purely pseudoscalar case, on the contrary, corrections are always negative. For $m_{S}<10\leavevmode\nobreak\ {\rm GeV}$, corrections are of the order of $-c_{t}^{2}\times 2\%$ for large invariant masses, growing in absolute value up to $-c_{t}^{2}\times 5\%$ at the threshold. Increasing $m_{S}$ the absolute value of the corrections decreases for large $m(t\bar{t})$ (halved for $m_{S}=300\leavevmode\nobreak\ {\rm GeV}$) while it grows at the threshold (doubled for $m_{S}=300\leavevmode\nobreak\ {\rm GeV}$). As already said, the maximally mixed case (right plot) shows distributions that correspond to the average of the analogous ones in the left and central plot. Clearly, by varying the angle $\phi$ that parameterises the mixing of the CP-even and CP-odd component, very different shapes can be obtained.
@@ 4.1 Statistical method, theory predictions and data
Our goal is to present an exploratory study for the potential bounds that can be obtained at the LHC on the $c_{t}$ (CP-even) and $\tilde{c}_{t}$ (CP-odd) couplings. We neither aim to fully exploit the potential of differential distributions that can be already now measured with high accuracy at the LHC, nor to quantify the projections for the full data-set of the High-Luminosity program; these aspects are left for future studies. Here, we want simply to show how the measurements of such process can lead to constraints for $c_{t}$ (CP-even) and $\tilde{c}_{t}$ (CP-odd) couplings of a top-philic additional scalar and the estimate the corresponding bounds.
@@
The theory prediction $\vec{\sigma}_{\rm th.}$ is given by $\sigma^{\rm mult.}_{\rm SM+NP}$, see the expression in Eq. (39), evaluated for the same binning of the data considered, $\vec{\sigma}_{\mathrm{obs.}}$. In fact, for our exploratory study we use pseudo-data for $\vec{\sigma}_{\mathrm{obs.}}$. First, we have considered several top-quark pair analyses ATLAS:2015lsn ; ATLAS:2016pal ; ATLAS:2019hxz ; CMS:2016oae ; CMS:2018adi ; CMS:2018htd in order to select a binning that is not only useful for our fit but also realistic. Given features discussed in Sec. 3, we have identified the CMS measurement in the lepton+jets channel of Ref. CMS:2018htd for the $m(t\bar{t})$ distribution as the most useful, since it has a large number of bins in the low $m(t\bar{t}$) region. Then, we have generated pseudo data assuming the SM only [^8] In this work, the only exception to this procedure is what is done in Sec. 7.2, as explained therein.: we have calculated for the binning of $m(t\bar{t})$ of Ref. CMS:2018htd $\sigma^{\rm mult.}_{\rm SM+NP}$ with $c_{t}=\tilde{c}_{t}=0$, in other words, the best SM best prediction, the quantity $\sigma^{\rm mult.}_{\rm SM}$ defined in Eq. (36). This procedure fully defines the quantities $\vec{\sigma}_{\mathrm{obs.}}$ and $\vec{\sigma}_{\rm th.}$. Further information entering the fit (binning, values of $\sigma^{\rm mult.}_{\rm SM+NP}$, etc.) is reported in Appendix D.
The last piece of information that has to be specified is the covariance matrix $\mathbf{V}^{-1}$. In Sec. 7 we will perform the same study discussed in this Section with the Higgs boson $H$ in the place of $S$. For consistency, $\vec{\sigma}_{\mathrm{obs.}}$ and $\mathbf{V}^{-1}$ will be the same used here and in particular it will be more clear the choice we have taken for $\mathbf{V}^{-1}$ and we explain in the following. Starting from the value of $\mathbf{V}^{-1}$ in Ref. CMS:2018htd , we have rescaled the entries of the matrix by a $(2.5)^{2}$ factor, [^9] This rescaling is equivalent to reduce by a factor 2.5 the experimental errors from Ref. CMS:2018htd . In practice, we have tuned this number in order to obtain what is described in the following sentence in the main text. such that in the case of the Higgs boson discussed in Sec. 7 we find for the CP-even component of top-Higgs interaction the same uncertainty reported in the aforecited Refs. CMS:2019art ; CMS:2020djy . In this way we can mimic, in a simplified statistical framework, the precision already achieved in the experimental analyses based on an integrated luminosity of $35.8\,\rm fb^{-1}$.